Would you rather catch an egg over a long period of time or short period of time? Explain your answer using impulse
1
Expert's answer
2019-10-15T14:01:53-0400

Let's write down the Newton's second law


"\vec F = {{d\vec p} \over {dt}}"

We can approximately write differentials as small differences (and let's work with magnitudes)


"\left| {\vec F} \right| \approx {{\left| {\Delta \vec p} \right|} \over {\Delta t}}"

The impulse change "{\left| {\Delta \vec p} \right|}" is fixed in our problem (it depends on the velocity of the egg and it's mass) because the final impulse of the egg (after catching) "\left| {\vec p} \right| = 0". We need to make "\left| {\vec F} \right|" as small as possible to protect the egg from breaking, so we should make "\Delta t" as big as possible. That means we should catch an egg over a long period of time.